Aqual:

x^2 + 10x + 16 = y find the vertex

MxxnLight:

@jhonyy9

jhonyy9:

so this is so easy - just you need get two numbers with sum equal 10 and product equal 16

Aqual:

huh?

jhonyy9:

what maybe these two numbers ?

Aqual:

what are h & k?

jhonyy9:

4*4 = 16 but 4+4=8

Aqual:

ohhh

Aqual:

ok

jhonyy9:

no bc. 4*8 = 32 not 16

jhonyy9:

and 4+8 = 12 not 10

Aqual:

1.6=x?

Aqual:

1.6*10=16

jhonyy9:

integers

Aqual:

oh

xxDeppressionxx:

Oh so you do 4=4=8•4=32

jhonyy9:

@supie

xxDeppressionxx:

4+4*

jhonyy9:

no a+b=10 a*b=16

Aqual:

2+8 = 10 2*8 16

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Aqual 2+8 = 10 2*8 16 \(\color{#0cbb34}{\text{End of Quote}}\) perfect

Aqual:

so how do i put that into vertex form

jhonyy9:

(x+8)(x+2)

xxDeppressionxx:

it is 8

Aqual:

so (8, 2)

jhonyy9:

yes x_1 = 8 x_2 = 2

jhonyy9:

@supie @Laylalyssa